# Can a MISL Fly? Analysis and Ingredients for Mutual Information Skill Learning

### v2

@@ 6.3 Ablation Studies
(2) Maximizing the information bottleneck is important. In Sec. 4.2, we interpret the intrinsic reward in METRA as a lower bound on an information bottleneck. We conduct ablation experiments to study the effect of maximizing this information bottleneck over maximizing the mutual information directly, a strategy typically used by prior methods (Eysenbach et al., 2019; Mendonca et al., 2021; Hansen et al., 2020). Results in Fig. 3 (Center) show that CSF failed to discover skills when only maximizing the mutual information (i.e. including the anti-exploration term). These results indicate that using the information bottleneck as the intrinsic reward may be important for MISL algorithms.
(3) Parameterization is key for CSF. When optimizing a lower bound on the mutual information $I^{\pi}(S,S^{\prime};Z)$ using a variational distribution, there are many ways to parametrize the critic $f(s,s^{\prime},z)$. In Eq. 9, we chose the parameterization $(\phi(s^{\prime})-\phi(s))^{\top}z$, but there are many other choices. Testing the sensitivity of this choice of parameterization allows us to determine whether a specific form of the lower bound is important. In Fig. 3, we study several variants of CSF that use (1) a monolithic network $\text{MLP}(s,s^{\prime})^{\top}z$ , (2) a Gaussian kernel ($-\frac{1}{2}||\phi(s^{\prime})-\phi(s)||_{2}^{2}$), or (3) a Laplacian kernel ($-||\phi(s^{\prime})-\phi(s)||_{1}$) as the critic parameterization. We find the alternative parameterizations are catastrophic for performance, suggesting that the inner product parameterization is key to CSF. We provide some insights for this parameterization in Appendix A.9.
@@ 6.4 CSF Matches SOTA for both Exploration and Downstream performance
Figure 4: CSF performs on par with METRA. We compare CSF with baselines on state coverage (left), zero-shot goal reaching (middle), and hierarchical control (right). CSF performs roughly on par with METRA and outperform all other baselines in most settings. Shaded areas indicate one standard deviation. Appendix Fig. 567 show the learning curves for all tasks.
Our final set of experiments compare CSF to prior MISL algorithms, measuring performance on both unsupervised exploration and solving downstream tasks.
@@
Taken together, CSF is a competitive MISL algorithm that matches the current SOTA. On the full set of results (Appendices B.3B.4, and B.5) we find that CSF continues to perform roughly on par with METRA on most tasks, though there are some tasks where CSF performs better and vice versa.
@@ 7 Conclusion
In this paper, we show our understanding of a current SOTA unsupervised skill discovery algorithms through the lens of MISL. Our analysis allowed the development of a simpler method CSF, which performs on par with METRA in most settings. More broadly, we provide evidence that mutual information maximization can still be effective to build high performing skill discovery algorithms.
@@ A.2 Proof of Proposition 1
where $10\leq\alpha<$ 1. Then, there exists a $11/\alpha>$ 1 that scales the expectation in Eq. 12 to exactly $1$:
@@
which contradicts the assumption that $\phi^{\star}$ is optimal. When $\mathbb{E}_{p(z)p^{\beta}(s,s^{\prime}\mid z)}\left[(\phi^{\star}(s^{\prime})-\phi^{\star}(s))^{\top}z\right]<0$, taking $-\phi^{\star}/\alpha$ gives us the same result. Therefore, we conclude that the optimal $\phi^{\star}$ must satisfy $\mathbb{E}_{p^{\beta}(s,s^{\prime})}\left[\lVert\phi^{\star}(s^{\prime})-\phi^{\star}(s)\rVert_{2}^{2}\right]=1$. ∎
@@ A.3 Proof of Proposition 2
[…]
=$\stackrel{{\scriptstyle\emph{(a)}}}{{=}}\log\frac{C_{d}(0)}{C_{d}(\lVert\phi(s^{\prime})-\phi(s)\rVert_{2})}$
@@
$\log\frac{\Gamma(d/2)2^{d/2-1}{\mathcal{I}}_{d/2-1}(\lVert\phi(s^{\prime})-\phi(s)\rVert_{2})}{\lVert\phi(s^{\prime})-\phi(s)\rVert_{2}^{d/2-1}}\\ =\frac{1}{2d}\lVert\phi(s^{\prime})-\phi(s)\rVert_{2}^{2}+O(\lVert\phi(s^{\prime})-\phi(s)\rVert_{2}^{3})\$
@@
$\lambda_{0}(d)(1-\lVert\phi(s^{\prime})-\phi(s)\rVert_{2}^{2})\approx-\log\mathbb{E}_{p}(z)\left[e^{(\phi(s^{\prime})-\phi(s))^{\top}z}\right]+\text{const.}.$. .
@@ A.4 Proof of Proposition 3
where $\text{LB}_{+}^{\pi}(\phi)$ is exactly the same as the RL objective $J(\pi)$ (Eq. 6). This lower bound is similar to Eq. 8 but it is under the target policy $\pi$ instead.
@@
=$\stackrel{{\scriptstyle\emph{(a)}}}{{=}}-\frac{1}{N}\sum_{i=1}^{N}\left[\log\left(\frac{1}{N}\sum_{j=1}^{N}C_{d}(\lVert\phi(s^{\prime}_{i})-\phi(s_{i})\rVert_{2})e^{\lVert\phi(s^{\prime}_{i})-\phi(s_{i})\rVert_{2}\frac{\left(\phi(s^{\prime}_{i})-\phi(s_{i})\right)^{\top}z_{j}}{\lVert\phi(s^{\prime}_{i})-\phi(s_{i})\rVert_{2}}}\right)\right.$
@@
=$\stackrel{{\scriptstyle\emph{(b)}}}{{=}}\hat{I}^{\pi}(S,S^{\prime};\phi(S^{\prime})-\phi(S))-\mathbb{E}_{p^{\pi}(s,s^{\prime})}\left[\log\frac{1}{C_{d}(\lVert\phi(s^{\prime}_{i})-\phi(s_{i})\rVert_{2})}\right]$
$\stackrel{{\scriptstyle\emph{(c)}}}{{\approx}}\hat{I}^{\pi}(S,S^{\prime};\phi(S^{\prime})-\phi(S))-\lambda_{0}(d)\mathbb{E}_{p^{\pi}(s,s^{\prime})}\left[\lVert\phi(s^{\prime})-\phi(s)\rVert_{2}^{2}\right]+\text{const.}$.
$\stackrel{{\scriptstyle\emph{(d)}}}{{\approx}}\hat{I}^{\pi}(S,S^{\prime};\phi(S^{\prime})-\phi(S))+\text{const.},$. ,
@@ A.5 Mutual Information Objectives used in Prior Methods
Prior MISL methods (Eysenbach et al., 2019; Gregor et al., 2016; Sharma et al., 2020; Hansen et al., 2020; Campos et al., 2020) adopt the min-max optimization procedure (Eq. 23) and use the same functional form of the lower bound on different variants of the mutual information as their objectives. We elaborate which mutual information each prior method optimizes next.
For DIAYN (Eysenbach et al., 2019) and VISR (Hansen et al., 2020), both representation learning and policy learning objectives are (up to some constants) lower bounds on $I(S;Z)\gtrsim-\mathbb{E}_{p(z)}[\log p(z)]+\mathbb{E}_{p(s,z)}[\log q(z\mid s)]$, where $\gtrsim$ denotes $>$ up to constant scaling or shifting. See Eq. 2 &\&& 3 of (Eysenbach et al., 2019) and Eq. 9 of (Hansen et al., 2020) for details.
@@
For DADS (Sharma et al., 2020), both representation learning and policy learning objectives are (approximate) lower bounds on $I(S^{\prime};Z\mid S)\geq-\mathbb{E}_{p(s^{\prime}\mid s)}[\log p(s^{\prime}\mid s)]+\mathbb{E}_{p(s^{\prime},z\mid s)}[\log q(s^{\prime}\mid s,z)]$. See Eq. 5 &\&& 6 of (Sharma et al., 2020) for details.
@@ A.6 The effect of the scaling coefficient $\xi$
=$\stackrel{{\scriptstyle\emph{(b)}}}{{=}}-\mathbb{E}_{p^{\beta}(s,s^{\prime})}\left[\log e^{0}\right]$
@@ A.8 Connection Between Representations Learned by METRA and Contrastive Representations
[…]
@@
[…]
We next propose a Lemma that relates a isotropic Gaussian distribution to a von Mises–Fisher distribution (Wikipedia, 2024) and then draw the connection between Claim 1 and Claim 2.
@@
[…]
@@
[…]
@@ A.9 Insights for the inner production critic parameterization
Our ablation studies in Sec. 6.3 shows that the inner product parameterization of the critic $f(s,s^{\prime},z)=(\phi(s^{\prime})-\phi(s))^{\top}$ is important for CSF (Fig. 3 (Right)). There are two explanations for these observations. First, the inner product parameterization tries to push together the difference of the representation of transitions $\phi(s^{\prime})-\phi(s)$ and the corresponding skill $z$, instead of focusing on representations of individual states $\phi(s)$ or ϕ(s)italic-$\phi(s^{\prime})$. This intuition is inline with the observation from prior work that mutual information $I(S;Z)$ is invariant to bijective mappings on $S$ (See Fig. 2 of (Park et al., 2024)), e.g., translation and scaling, indicating that maximizing the mutual information between change of states and skills ($I(S,S^{\prime};Z)$) encourages better state space coverage. Second, the inner product parameterization allows us to analytically compute the second-order Taylor approximation in Proposition 2, drawing the connection between the METRA representation objective and contrastive learning.
@@ A.10 Intuition for zero-shot goal inference
In zero-shot goal reaching setting, we want to figure out the corresponding latent skill $z$ when given a goal state or image $g$; a problem which can be cast as inferring the $z\in{\mathcal{Z}}$ that maximizes the posterior $p^{\pi}(z\mid s,g)$. Since the ground truth posterior $p^{\pi}(z\mid s,g)$ is unknown, a typical workaround is first estimating a variational approximation of $p^{\pi}(z\mid s,g)$ and then maximizing the variational posterior $q(z\mid s,g)$. We provide an intuition for zero-shot goal inference by specifying the variational posterior as
@@ B.1 Simplicity of CSF compared to METRA
On the one hand, since our algorithm prevents the dual gradient descent optimization in the METRA representation objective, we do not have the slack variable $\epsilon$, do not have to specify the initial value of the dual variable $\lambda$, and remove the dual variable optimizer with its learning rate, resulting in three less hyperparameters. On the other hand, we introduce one coefficient $\xi$ to scale the negative term $\text{LB}^{\beta}_{-}(\phi)$ in the contrastive lower bound, which has an effect similar to the tradeoff coefficient used in information bottleneck optimization ($\beta$ in (Alemi et al., 2017)). Taken together, CSF uses three less hyperparameters than METRA and introduces one hyperparameter to balance the positive term ($\text{LB}^{\beta}_{+}(\phi)$) and the negative term ($\text{LB}^{\beta}_{-}(\phi)$) in the representation objective.
@@ B.3 Exploration performance
For CSF, all tasks were trained with continuous $z$ sampled from a uniform vMF distribution and $\lambda=5$. METRA also uses a continuous $z$ sampled from a uniform vMF distribution for all environments except for HalfCheetah and Kitchen, where we used a one-hot discrete $z$, consistent with the original work (Park et al., 2024). CIC uses a continuous $z$ sampled from a standard Gaussian for all environments. DIAYN uses a one-hot discrete $z$ for all environments. DADS uses a continuous $z$ sampled from a uniform distribution on $[-1,1]$ for all environments. Finally, VISR uses a continuous $z$ sampled from a uniform vMF distribution for all environments. Please refer to Table 2 for a full overview of skill dimensions per method and environment. A table with all relevant hyperparameters for the unsupervised training phase can be found in Table 1.
@@
Figure 5: State space coverage. We plot the unique number of coordinates visited by the agent, except for Kitchen where we plot the task coverage. We find CSF matches the prior state-of-the-art MISL algorithms on $4/6$ tasks, and strongly outperforms METRA in Robobin. Shaded areas indicate one standard deviation.

### v3

@@ 6.3 Ablation Studies
(2) Maximizing the information bottleneck is important. In Sec. 4.2, we interpret the intrinsic reward in METRA as a lower bound on an information bottleneck. We conduct ablation experiments to study the effect of maximizing this information bottleneck over maximizing the mutual information directly, a strategy typically used by prior methods (Eysenbach et al., 2019; Mendonca et al., 2021; Hansen et al., 2020). Results in Fig. 3 (Center) show that CSF failed to discover skills when only maximizing the mutual information (i.e., including the anti-exploration term). These results indicate that using the information bottleneck as the intrinsic reward may be important for MISL algorithms.
[…]
@@ 6.4 CSF Matches SOTA for both Exploration and Downstream performance
Figure 4: CSF performs on par with METRA. We compare CSF with baselines on state coverage (left), zero-shot goal reaching (middle), and hierarchical control (right). CSF performs roughly on par with METRA and outperforms all other baselines in most settings. Shaded areas indicate one standard deviation. Appendix Fig. 567 show the learning curves for all tasks.
Our final set of experiments compares CSF to prior MISL algorithms, measuring performance on both unsupervised exploration and solving downstream tasks.
@@
Taken together, CSF is a competitive MISL algorithm that matches the current SOTA. On the full set of results (Appendices B.3B.4, and B.5), we find that CSF continues to perform roughly on par with METRA on most tasks, though there are some tasks where CSF performs better and vice versa.
@@ 7 Conclusion
In this paper, we show our understanding of a current SOTA unsupervised skill discovery algorithm through the lens of MISL. Our analysis allowed the development of a simpler method CSF, which performs on par with METRA in most settings. More broadly, we provide evidence that mutual information maximization can still be effective in building high-performing skill discovery algorithms.
@@ A.2 Proof of Proposition 1
where $0\leq\alpha<1$. Then, there exists a $1/\alpha>1$ that scales the expectation in Eq. 12 to exactly $1$:
@@
which contradicts the assumption that $\phi^{\star}$ is optimal. When $\mathbb{E}_{p(z)p^{\beta}(s,s^{\prime}\mid z)}\left[(\phi^{\star}(s^{\prime})-\phi^{\star}(s))^{\top}z\right]<0$, taking $\phi^{\star}/\alpha$ gives us the same result. Therefore, we conclude that the optimal $\phi^{\star}$ must satisfy $\mathbb{E}_{p^{\beta}(s,s^{\prime})}\left[\lVert\phi^{\star}(s^{\prime})-\phi^{\star}(s)\rVert_{2}^{2}\right]=1$. ∎
@@ A.3 Proof of Proposition 2
$\hskip 10.00002pt+\log\int C_{d}(\lVert\phi(s^{\prime})-\phi(s)\rVert_{2})e^{\lVert\phi(s^{\prime})-\phi(s)\rVert_{2}\frac{(\phi(s^{\prime})-\phi(s))^{\top}z}{\lVert\phi(s^{\prime})-\phi(s)\rVert_{2}}}dz$
$\stackrel{{\scriptstyle\emph{(a)}}}{{=}}\log\frac{C_{d}(0)}{C_{d}(\lVert\phi(s^{\prime})-\phi(s)\rVert_{2})}$
@@
$\log\frac{\Gamma(d/2)2^{d/2-1}{\mathcal{I}}_{d/2-1}(\lVert\phi(s^{\prime})-\phi(s)\rVert_{2})}{\lVert\phi(s^{\prime})-\phi(s)\rVert_{2}^{d/2-1}}\\ =\frac{1}{2d}\lVert\phi(s^{\prime})-\phi(s)\rVert_{2}^{2}+O(\lVert\phi(s^{\prime})-\phi(s)\rVert_{2}^{3})\\$
@@
$\lambda_{0}(d)(1-\lVert\phi(s^{\prime})-\phi(s)\rVert_{2}^{2})\approx-\log\mathbb{E}_{p}(z)\left[e^{(\phi(s^{\prime})-\phi(s))^{\top}z}\right]+\text{const.}.$
@@ A.4 Proof of Proposition 3
where $\text{LB}_{+}^{\pi}(\phi)$ is exactly the same as the RL objective $J(\pi)$ (Eq. 6). This lower bound is similar to Eq. 8, but it is under the target policy $\pi$ instead.
@@
$\stackrel{{\scriptstyle\emph{(a)}}}{{=}}-\frac{1}{N}\sum_{i=1}^{N}\left[\log\left(\frac{1}{N}\sum_{j=1}^{N}C_{d}(\lVert\phi(s^{\prime}_{i})-\phi(s_{i})\rVert_{2})e^{\lVert\phi(s^{\prime}_{i})-\phi(s_{i})\rVert_{2}\frac{\left(\phi(s^{\prime}_{i})-\phi(s_{i})\right)^{\top}z_{j}}{\lVert\phi(s^{\prime}_{i})-\phi(s_{i})\rVert_{2}}}\right)\right$.
@@
$\stackrel{{\scriptstyle\emph{(b)}}}{{=}}\hat{I}^{\pi}(S,S^{\prime};\phi(S^{\prime})-\phi(S))-\mathbb{E}_{p^{\pi}(s,s^{\prime})}\left[\log\frac{1}{C_{d}(\lVert\phi(s^{\prime}_{i})-\phi(s_{i})\rVert_{2})}\right]$
$\stackrel{{\scriptstyle\emph{(c)}}}{{\approx}}\hat{I}^{\pi}(S,S^{\prime};\phi(S^{\prime})-\phi(S))-\lambda_{0}(d)\mathbb{E}_{p^{\pi}(s,s^{\prime})}\left[\lVert\phi(s^{\prime})-\phi(s)\rVert_{2}^{2}\right]+\text{const.}$
$\stackrel{{\scriptstyle\emph{(d)}}}{{\approx}}\hat{I}^{\pi}(S,S^{\prime};\phi(S^{\prime})-\phi(S))+\text{const.},$
@@ A.5 Mutual Information Objectives used in Prior Methods
Prior MISL methods (Eysenbach et al., 2019; Gregor et al., 2016; Sharma et al., 2020; Hansen et al., 2020; Campos et al., 2020) adopt the min-max optimization procedure (Eq. 23) and use the same functional form of the lower bound on different variants of the mutual information as their objectives. We elaborate on which mutual information each prior method optimizes next.
For DIAYN (Eysenbach et al., 2019) and VISR (Hansen et al., 2020), both representation learning and policy learning objectives are (up to some constants) lower bounds on $I(S;Z)\gtrsim-\mathbb{E}_{p(z)}[\log p(z)]+\mathbb{E}_{p(s,z)}[\log q(z\mid s)]$, where $\gtrsim$ denotes >> up to constant scaling or shifting. See Eq. 2 &\& 3 of (Eysenbach et al., 2019) and Eq. 9 of (Hansen et al., 2020) for details.
@@
For DADS (Sharma et al., 2020), both representation learning and policy learning objectives are (approximate) lower bounds on $I(S^{\prime};Z\mid S)\geq-\mathbb{E}_{p(s^{\prime}\mid s)}[\log p(s^{\prime}\mid s)]+\mathbb{E}_{p(s^{\prime},z\mid s)}[\log q(s^{\prime}\mid s,z)]$. See Eq. 5 &\& 6 of (Sharma et al., 2020) for details.
@@ A.6 The effect of the scaling coefficient $\xi$
$\stackrel{{\scriptstyle\emph{(b)}}}{{=}}-\mathbb{E}_{p^{\beta}(s,s^{\prime})}\left[\log e^{0}\right]$
@@ A.8 Connection Between Representations Learned by METRA and Contrastive Representations
The state representations $\phi$ learned by METRA satisfies that $\phi(s^{\prime})-\phi(s)-z\xrightarrow{\text{d}}{\mathcal{N}}(0,\sigma_{\phi}^{2}I)$, or, equivalently, $\phi(s^{\prime})-\phi(s)\mid z\xrightarrow{\text{d}}{\mathcal{N}}(z,\sigma_{\phi}^{2}I)$, where $\xrightarrow{\text{d}}$ denotes convergence in distribution and $\sigma_{\phi}$ is the standard deviation of the isotropic Gaussian.
@@
The state representations $\phi$ learned by METRA also satisfy $\frac{\phi(s^{\prime})-\phi(s)}{\lVert\phi(s^{\prime})-\phi(s)\rVert_{2}}\xrightarrow{\text{d}}\textsc{Unif}({\mathbb{S}}^{d-1})$.
We next propose a Lemma that relates an isotropic Gaussian distribution to a von Mises–Fisher distribution (Wikipedia, 2024) and then draw the connection between Claim 1 and Claim 2.
@@
Since the didactic experiments in Sec. 6.2 have shown that $\phi(s^{\prime})-\phi(s)\mid z$ converges to a Gaussian distribution ${\mathcal{N}}(z,\sigma_{\phi}^{2}I)$ (Claim 1) and note that $\lVert z\rVert_{2}=1$, by applying Lemma A.8, we conjecture that restricting $\phi(s^{\prime})-\phi(s)$ within $\{\lVert\phi(s^{\prime})-\phi(s)\rVert_{2}=1\}$ produces a von Mises-Fisher distribution, i.e., $\left.\frac{\phi(s^{\prime})-\phi(s)}{\lVert\phi(s^{\prime})-\phi(s)\rVert_{2}}\right\rvert z\xrightarrow{d}\textsc{vMF}(z,1/\sigma_{\phi}^{2})$. Furthermore, we can derive the marginal density of $\frac{\phi(s^{\prime})-\phi(s)}{\lVert\phi(s^{\prime})-\phi(s)\rVert_{2}}$,
@@
$\stackrel{{\scriptstyle\emph{(a)}}}{{=}}C_{d}(0)\int C_{d}\left(\frac{1}{\sigma_{\phi}^{2}}\right)\exp\left(\frac{1}{\sigma_{\phi}^{2}}\cdot\frac{(\phi(s^{\prime})-\phi(s))^{\top}z}{\lVert\phi(s^{\prime})-\phi(s)\rVert_{2}}\right)dz$
@@ A.9 Insights for the inner production critic parameterization
Our ablation studies in Sec. 6.3 shows that the inner product parameterization of the critic $f(s,s^{\prime},z)=(\phi(s^{\prime})-\phi(s))^{\top}$ is important for CSF (Fig. 3 (Right)). There are two explanations for these observations. First, the inner product parameterization tries to push together the difference of the representation of transitions $\phi(s^{\prime})-\phi(s)$ and the corresponding skill $z$, instead of focusing on representations of individual states $\phi(s)$ or $\phi(s^{\prime})$. This intuition is inline with the observation from prior work that mutual information $I(S;Z)$ is invariant to bijective mappings on $S$ (See Fig. 2 of (Park et al., 2024)), e.g., translation and scaling, indicating that maximizing the mutual information between change of states and skills ($I(S,S^{\prime};Z)$) encourages better state space coverage. Second, the inner product parameterization allows us to analytically compute the second-order Taylor approximation in Proposition 2, drawing the connection between the METRA representation objective and contrastive learning.
@@ A.10 Intuition for zero-shot goal inference
In the zero-shot goal-reaching setting, we want to figure out the corresponding latent skill $z$ when given a goal state or image $g$; a problem which can be cast as inferring the $z\in{\mathcal{Z}}$ that maximizes the posterior $p^{\pi}(z\mid s,g)$. Since the ground truth posterior $p^{\pi}(z\mid s,g)$ is unknown, a typical workaround is first estimating a variational approximation of $p^{\pi}(z\mid s,g)$ and then maximizing the variational posterior $q(z\mid s,g)$. We provide an intuition for zero-shot goal inference by specifying the variational posterior as
@@ B.1 Simplicity of CSF compared to METRA
On the one hand, since our algorithm prevents the dual gradient descent optimization in the METRA representation objective, we do not have the slack variable $\epsilon$, do not have to specify the initial value of the dual variable $\lambda$, and remove the dual variable optimizer with its learning rate, resulting in three fewer hyperparameters. On the other hand, we introduce one coefficient $\xi$ to scale the negative term $\text{LB}^{\beta}_{-}(\phi)$ in the contrastive lower bound, which has an effect similar to the tradeoff coefficient used in information bottleneck optimization ($\beta$ in (Alemi et al., 2017)). Taken together, CSF uses three less hyperparameters than METRA and introduces one hyperparameter to balance the positive term ($\text{LB}^{\beta}_{+}(\phi)$) and the negative term ($\text{LB}^{\beta}_{-}(\phi)$) in the representation objective.
@@ B.3 Exploration performance
For CSF, all tasks were trained with continuous $z$ sampled from a uniform vMF distribution and $\lambda=5$. METRA also uses a continuous $z$ sampled from a uniform vMF distribution for all environments except for HalfCheetah and Kitchen, where we used a one-hot discrete $z$, consistent with the original work (Park et al., 2024). CIC uses a continuous $z$ sampled from a standard Gaussian for all environments. DIAYN uses a one-hot discrete $z$ for all environments. DADS uses a continuous $z$ sampled from a uniform distribution on [1,1][-1,1] for all environments. Finally, VISR uses a continuous $z$ sampled from a uniform vMF distribution for all environments. Please refer to Table 2 for a full overview of skill dimensions per method and environment. A table with all relevant hyperparameters for the unsupervised training phase can be found in Table 1.
@@
Figure 5: State space coverage. We plot the unique number of coordinates visited by the agent, except for the Kitchen where we plot the task coverage. We find CSF matches the prior state-of-the-art MISL algorithms on $4/6$ tasks, and strongly outperforms METRA in Robobin. Shaded areas indicate one standard deviation.